# A novel characterization of structures in smooth regression curves: from a viewpoint of persistent homology

### v3

@@ 5.3 Modality
We provide an illustrative example to infer the modality of a smooth regression function from noisy samples using estimated barcodes of its first derivative. We characterize modes of a smooth regression function $f$ on [-1, 1] using the critical points of $f$. We say that a critical point $c\in$ [-1, 1] is a mode of $f$ if there exists a $\delta>0$ such that $f$ is concave in [$c$ $\delta$, $c$ + $\delta$]. We use the method of zero crossings of the derivative to estimate the critical points.
We simulate noisy observations from the mixture of two normal distributions truncated on [-1, 1] with mean -0.5 and 0.5, with the same variance 0.1, and equal mixture proportions. The graph of true and estimated function is given in Figure 13 along with the SiZer map of the data. We obtain critical points of the estimated function using significant zero crossings of the derivative. We focus on two of the critical points depicted by blue dots in Figure 13, and aim to infer the modality of the underlying true function in the local regions highlighted by green vertical dotted lines in Figure 13. These regions are $\delta$-neighbourhood around the critical points of the estimated function, where we choose $\delta=0.1$. We use the aforementioned characterization to infer whether the critical points represent modes of the true function.
@@
We obtain barcodes of the estimated derivative using the procedure given in Section 3, and infer the monotonicity of the function in the local regions $[c-\delta,c]$ and $[c,c+\delta]$ to infer modality. Note that, here $c$ represents the critical points of interest, and $\delta=0.1$. We denote the critical points of the estimated function $\hat{m}$ depicted in Figure 13 by $\boldsymbol{\hat{c}_{1}}$ and $\boldsymbol{\hat{c}_{2}}$. The estimated barcodes for the Gaussian kernel are shown in Figure 14 for the sample size $n=50$, and the bandwidth $h=n^{-1/3}$. The estimated critical values $\boldsymbol{\hat{c}_{1}}$ and $\boldsymbol{\hat{c}_{2}}$ are -0.47, 0.51, respectively. The estimated barcodes in the local regions $[\boldsymbol{\hat{c}_{1}}-\delta,\boldsymbol{\hat{c}_{1}}],[\boldsymbol{\hat{c}_{1}},\boldsymbol{\hat{c}_{1}}+\delta],[\boldsymbol{\hat{c}_{2}}-\delta,\boldsymbol{\hat{c}_{2}}],[\boldsymbol{\hat{c}_{2}},\boldsymbol{\hat{c}_{2}}+\delta]$ are $\boldsymbol{\{[0.05,0.72]\},\{[-0.72,0.03]\},\{[0.09,0.81]\},\{[-0.69,0.07]\}}$, respectively, for $\delta=0.1$. Note that, we highlight the true and estimated death or birth times in different regions using black and green dotted lines, respectively, to investigate monotonicity in these regions (see, Section 5.1). It is evident from the estimated barcodes of $\hat{m}_{1}$ that $\hat{m}$ is increasing on $[\boldsymbol{\hat{c}_{1}}-0.1,\boldsymbol{\hat{c}_{1}}]$, since the death time (0.05) of the observed feature is positive. On the other hand, the birth time of the feature in the estimated barcode on $[\boldsymbol{\hat{c}_{1}},\boldsymbol{\hat{c}_{1}}+0.1]$ is non-negative yielding that $\hat{m}$ is not non-increasing on $[\boldsymbol{\hat{c}_{1}},\boldsymbol{\hat{c}_{1}}+0.1]$. Hence, we conclude that the critical point $\boldsymbol{\hat{c}_{1}}$ is not a mode of the estimated function. Similar conclusion also holds for $\boldsymbol{\hat{c}_{2}}$. It can be seen from Table 7 that the observed critical points are not significant. This fact can be validated by the SiZer map of the data in Figure 13. Since for $h=n^{-1/3}$ (black dotted line) there is a purple region between the two blue and red regions indicating that the observed critical points are not significant.
Figure 14: Estimated barcodes of $\hat{m}_{1}$ for the Gaussian kernel for $h=n^{-1/3},n=50$.
Table 7: Significance of modes for Gaussian Kernel.
Barcode for $\boldsymbol{\hat{c}_{1}}$ Barcode for $\boldsymbol{\hat{c}_{2}}$
On $[\mathbf{\hat{c}_{1}-0.1},\mathbf{\hat{c}_{1}}]$ On $[\mathbf{\hat{c}_{1}},\mathbf{\hat{c}_{1}+0.1}]$ On $[\mathbf{\hat{c}_{2}-0.1},\mathbf{\hat{c}_{2}}]$ On $[\mathbf{\hat{c}_{2}},\mathbf{\hat{c}_{2}+0.1}]$
$\mathbf{a_{1}}$ $\mathbf{b_{1}}$ $\mathbf{a_{1}}$ $\mathbf{b_{1}}$ $\mathbf{a_{1}}$ $\mathbf{b_{1}}$ $\mathbf{a_{1}}$ $\mathbf{b_{1}}$
$\mathbf{\hat{c}_{p}}$0.46 -0.56 -0.36 -0.46 0.51 0.41 0.61 0.51
$\boldsymbol{\hat{\beta}_{1}}$0.27 2.6 -2.7 -0.39 -0.05 2.3 -2.2 -0.16
$\mathbf{p}$-value 0.84 0.05 0.02 0.7 0.9 0.06 0.04 0.8
Next, we obtain the barcodes of $\hat{m}_{1}$ using the Cauchy kernel. The critical values to be investigated to identify modes are $\boldsymbol{\hat{c}_{1}=-0.51}$ and $\boldsymbol{\hat{c}_{2}=0.51}$. The estimated barcodes in the local regions $[\boldsymbol{\hat{c}_{1}}-0.1,\boldsymbol{\hat{c}_{1}}],[\boldsymbol{\hat{c}_{1}},\boldsymbol{\hat{c}_{1}}+0.1],[\boldsymbol{\hat{c}_{2}}-0.1,\boldsymbol{\hat{c}_{2}}],[\boldsymbol{\hat{c}_{2}},\boldsymbol{\hat{c}_{2}}+0.1]$ are $\boldsymbol{\{[0.3,1.1]\}}$, $\boldsymbol{\{[-0.7,0.3]\},\{[0.01,0.91]\},\{[-0.94,-0.02]\}}$, respectively. These barcodes are shown in Figure 15. It is evident from the estimated barcode that only $\boldsymbol{\hat{c}_{2}}$ can be characterized to be a mode. However, it is not significant as can be seen in Table 8.
Figure 15: Estimated barcodes of $\hat{m}_{1}$ for the Cauchy kernel for $h=n^{-1/3},n=50$.
Table 8: Significance of modes for Cauchy Kernel.
Barcode for $\boldsymbol{\hat{c}_{1}}$ Barcode for $\boldsymbol{\hat{c}_{2}}$
On $[\mathbf{\hat{c}_{1}-0.1},\mathbf{\hat{c}_{1}}]$ On $[\mathbf{\hat{c}_{1}},\mathbf{\hat{c}_{1}+0.1}]$ On $[\mathbf{\hat{c}_{2}-0.1},\mathbf{\hat{c}_{2}}]$ On $[\mathbf{\hat{c}_{2}},\mathbf{\hat{c}_{2}+0.1}]$
$\mathbf{a_{1}}$ $\mathbf{b_{1}}$ $\mathbf{a_{1}}$ $\mathbf{b_{1}}$ $\mathbf{a_{1}}$ $\mathbf{b_{1}}$ $\mathbf{a_{1}}$ $\mathbf{b_{1}}$
$\mathbf{\hat{c}_{p}}$0.51 -0.59 -0.41 -0.5 0.51 0.41 0.61 0.51
$\boldsymbol{\hat{\beta}_{1}}$ 1.4 4.7 -3.4 1.006 -0.47 3.3 -3.8 -0.6
$\mathbf{p}$-value 0.3 0.0007 0.01 0.5 0.7 0.01 0.002 0.6
@@ 6 Real data study
We apply the proposed methodology to two real data sets to infer local structures such as monotonicity, convexity, and modality. We provide statistical significance of observed structures using the methodology proposed in Section 4, and compare the results with the respective SiZer maps. We investigate local structures in the first data set for two different bandwidths $h_{1}$, $h_{2}$, highlighted by black and gray horizontal dotted lines on the SiZer maps, respectively. On the other hand, we choose a bandwidth to observe a certain structure and investigate its significance for the second data set. In addition, we also focus on certain structures arising in certain local regions highlighted by green dotted lines in Figure 20. In this section, we use the Gaussian kernel supported on [-1, 1] to estimate the underlying true regression function.
@@ 6.1 Data set 1
We consider the intercountry lifecycle savings data available in the R software library, “datasets.” This data is based on the life-cycle savings hypothesis developed by Franco Modigliani (1975). According to the lifecycle savings hypothesis, the savings ratio (Y) (aggregate personal savings divided by disposable income) is explained by the following four variables: the percentage of the population less than 15 years old ($X_{1}$), the percentage of the population over 75 years old ($X_{2}$), the disposable income per capita ($X_{3}$), and the percentage rate of change in per-capita disposable income ($X_{4}$). This data set consists of observations on these four variables for 50 countries. We fit a non-parametric regression model (see, Equation (3.1)) for each of the four regressors $X_{1},X_{2},X_{3},X_{4}$ to infer local structures. A non-parametric regression fit ($\hat{m}$) for $X_{1}$, the estimated derivative ($\hat{m}_{1}$), SiZer map for $\hat{m}$, and the estimated barcodes of $\hat{m}_{1}$ for two different bandwidths, $h_{1}=15$, $h_{2}=8$, is shown in Figure 16. It is evident from the barcode of $\hat{m}_{1}$ that $\hat{m}$ is significantly decreasing for $h_{1}$, and (see, Figure 16). It can be seen in Table 9 that the observed monotone structure is significant at 5 % level of significance. Since the birth times of the 0-dimensional features in the estimated barcode of $\hat{m_{1}}$ is negative for $h_{1}$ (see, Figure 16) and significant (see, Table 9). However, for $h_{2}=8$, the birth times are negative but not significant implying that the observed structure is not significantly decreasing.
Figure 16: Comparison of local structures in the estimated regression function ($\hat{m}$) derived from the estimated derivative ($\hat{m}_{1}$) with the SiZer map for $X_{1}$ using the bandwidths $h_{1}=15$, $h_{2}=8$.
Now, we compare the inferred monotone structure from the barcode of the estimated derivative with the SiZer map of the data. Note that, for $h_{1}=15$ denoted by the black dotted line on the SiZer map in Figure 16, the observed structure is significantly decreasing as highlighted by the red region. The significance of features articulated in Table 9 suggests the same. Since the critical points, denoted by the green dots on the SiZer map, corresponding to the observed features lie on the red region on the SiZer map. This implies that function is significantly decreasing at these points. This further implies that $\hat{m}_{1}$ at these points is significantly negative which is consistent with the results in Table 9. However, for $h_{2}=8$, denoted by the gray dotted line, the SiZer map indicates the purple region for some x-locations implying that the observed structure is not significantly decreasing. It can be seen in Table 9 that the observed birth times ($b_{1}$) are not significant indicating that $\hat{m}$ can not be concluded to be significantly decreasing.
Table 9: Significance of local structures for $X_{1}$.
$\mathbf{h_{1}=15}$ $\mathbf{h_{2}=8}$
$\mathbf{a_{1}}$ $\mathbf{b_{1}}$ $\mathbf{a_{2}}$ $\mathbf{b_{2}}$ $\mathbf{a_{1}}$ $\mathbf{b_{1}}$ $\mathbf{a_{2}}$ $\mathbf{b_{2}}$
$\mathbf{\hat{c}_{p}}$ 35.8 47.1 26.7 45.5 35.8 46.5 38.01 22.5
$\boldsymbol{\hat{\beta}_{1}}$$2.7\times 10^{3}$3595.5 -3620.7 -3452.3 -926.8 1195.1 -854.8 946.3
$\mathbf{p}$-value 0.0005 0.0194 $\mathbf{0.0006}$ 0.006 0.0003 0.05 0.001 0.05
Now, for the regressors $X_{2}$, $X_{3}$ and $X_{4}$, some key insights about the observed structure in $\hat{m}$ are summarized as follows. The observed structure in $\hat{m}$ is not significantly monotone for each of the bandwidths for $X_{2}$ as suggested by the estimated barcodes in Figure 17 and Table 10. It can be seen in Table 10 that the observed features for $h_{1}$ are not significant while for $h_{2}$ the feature $[a_{1},b_{1}]$ is significant. Note that, for $h_{2}$, $a_{1}<0$ and $b_{1}>0$ indicating that $\hat{m}$ can not be a monotone function that is also evident from the SiZer map in Figure 17. However, for $h_{1}$ the SiZer map suggests significantly increasing structure contrary to the structure derived from the estimate barcode in Table 10.
The observed barcode in Figure 18 indicates that $\hat{m}$ is monotone increasing for the bandwidths $h_{1}=Q_{X}(0.95)$, and $h_{2}=Q_{X}(0.75)$, where $Q_{X}(u),u\in(0,1)$, denote the $u$-th empirical quantile of the underlying regressor $X$ (in this case $X_{3}$). However, the observed monotone structure is not significant for $h_{1}$ whereas significant for $h_{2}$ as evident from Table 11. In contrast, the SiZer map in Figure 18 indicates that the observed structure is not significantly monotone increasing.
The observed barcode in Figure 19 indicates that $\hat{m}$ is monotone increasing for the bandwidths $h_{1}=25$, and $h_{2}=Q_{X}(0.95)$. However, the observed monotone structure is significant only for the bandwidth $h_{2}=Q_{X}(0.95)$. In contrast, the SiZer map indicates that the observed monotone structure is significant for $h_{1}$, and insignificant for $h_{2}$.
Figure 17: Comparison of local structures in the estimated regression function ($\hat{m}$) derived from the estimated derivative ($\hat{m}_{1}$) with the SiZer map for $X_{2}$ using the bandwidths $h_{1}=4$, $h_{2}=0.5$.
Table 10: Significance of local structures for $X_{2}$.
$\mathbf{h_{1}=4}$ $\mathbf{h_{2}=0.5}$
$\mathbf{a_{1}}$ $\mathbf{b_{1}}$ $\mathbf{a_{1}}$ $\mathbf{b_{1}}$ $\mathbf{a_{2}}$ $\mathbf{b_{2}}$ $\mathbf{a_{3}}$ $\mathbf{b_{3}}$
$\mathbf{\hat{c}_{p}}$ 0.56 3.6 0.56 1.9 4.02 3.6 0.72 4.2
$\boldsymbol{\hat{\beta}_{1}}$ 1219.31 951.6 -58.4 28.4 -7.2 1.8 -64.09 -15.01
$\mathbf{p}$-value 0.01 0.02 $\mathbf{0.0003}$ 0.006 0.51 0.8 0.0009 0.21
Figure 18: Comparison of local structures in the estimated regression function ($\hat{m}$) derived from the estimated derivative ($\hat{m}_{1}$) with the SiZer map for $X_{3}$ using the bandwidths $h_{1}=Q_{X}(0.95)$, $h_{2}=Q_{X}(0.75)$.
Table 11: Significance of local structures for $X_{3}$ and $X_{4}$
Variable $X_{3}$ Variable $X_{4}$
$\mathbf{h_{1}=Q_{X}(0.95)}$ $\mathbf{h_{2}=Q_{X}(0.75)}$ $\mathbf{h_{1}=25}$ $\mathbf{h_{2}=Q_{X}(0.95)}$
$\mathbf{a_{1}}$ $\mathbf{b_{1}}$ $\mathbf{a_{1}}$ $\mathbf{b_{1}}$ $\mathbf{a_{1}}$ $\mathbf{b_{1}}$ $\mathbf{a_{1}}$ $\mathbf{b_{1}}$
$\mathbf{\hat{c}_{p}}$ 4001.8 1286.7 4001.8 568.07 0.22 12.6 16.7 11.3
$\boldsymbol{\hat{\beta}_{1}}$$2.7\times 10^{5}$ $4.8\times 10^{5}$$2.9\times 10^{5}$ $3.6\times 10^{5}$ 18652.5 14418.6 -1977.6 544.5
$\mathbf{p}$-value 0.6 0.08 $\mathbf{0.009}$ 0.01 0.01 0.04 0.003 0.56
Figure 19: Comparison of local structures in the estimated regression function ($\hat{m}$) derived from the estimated derivative ($\hat{m}_{1}$) with the SiZer map for $X_{4}$ using the bandwidths $h_{1}=25$, $h_{2}=Q_{X}(0.95)$.
@@ 6.2 Data set 2
We consider the “Motorcycle data” available in the R package “adlift”. This data set contains $n=133$ observations of simulations measuring the effects of motorcycle crashes on victims heads. There are two variables in the data set; time after a simulated impact with motorcycles (X) and head acceleration of a post-mortem human test object (Y). The observed structure in the data set is shown in Figure 20. We fit a non-parametric regression model (see, Equation 3.1) to the observed data, and aim to infer the observed structure using the proposed methodology. We compute the barcode of the estimated derivative using the procedure given in Section 3 to investigate if the observed structure is significant or not. In addition, we aim to investigate if the valley structure arising in the region [15, 25], and the observed peak in the the region [25, 35] is significant or not. These local regions are highlighted via green vertical dotted lines in Figure 20.
A non-parametric regression fit ($\hat{m}$), the estimated derivative ($\hat{m}_{1}$), the SiZer map for $\hat{m}$, and the estimated barcodes of $\hat{m}_{1}$ is shown in Figure 20. We choose the bandwidth $h=2.5$ highlighted by the black dotted line on the SiZer map, to estimate the regression function $m$. It is evident from Table 12 that none of the observed features in the barcode of $\hat{m}_{1}$ are significant. Hence, the observed structure in the data is not significant, which is also evident from the SiZer map in Figure 20. Furthermore, the observed valley in [15, 25] and the observed peak in [25, 35] are not significant, since from Table 13 none of the features in the observed barcode in [15, 25] and [25, 35] are significant. Note that, the same inference is also drawn from the SiZer map as there is a purple region between the red and blue region on [15, 25], and a purple region after the blue region on [25, 35]. This suggests that the observed peak and valley in the region [25, 35] and [15, 25], respectively, is not significant.
@@
Figure 20: Comparison of local structures in the estimated regression function ($\hat{m}$) derived from the estimated derivative ($\hat{m}_{1}$) with the SiZer map using the bandwidths $h=2.5$
Table 12: Significance of the observed local structures for $h=2.5$
$\mathbf{a_{1}}$ $\mathbf{b_{1}}$ $\mathbf{a_{2}}$ $\mathbf{b_{2}}$ $\mathbf{a_{3}}$ $\mathbf{b_{3}}$ $\mathbf{a_{4}}$ $\mathbf{b_{4}}$
$\mathbf{\hat{c}_{p}}$ 18.1 23.8 18.1 55.3 10.2 40.7 35.06 40.7
$\boldsymbol{\hat{\beta}_{1}}$2485.8 3434.2 -2485.8 -236.1 -578.6 -238.01 -1438.9 -238.01
$\mathbf{p}$-value $\mathbf{0.0305}$ 0.0001 $\mathbf{0.0305}$ 0.675 0.468 0.733 0.119 0.733
Table 13: Significance of convexity on [15, 25], and modality on [25, 35].
Convexity on [15, 25] Modality on [25, 35]
$\mathbf{a_{1}}$ $\mathbf{b_{1}}$ $\mathbf{a_{2}}$ $\mathbf{b_{2}}$ $\mathbf{a_{1}}$ $\mathbf{b_{1}}$
$\mathbf{\hat{c}_{p}}$ 18.1 23.8 18.1 17.04 10.2 23.8
$\boldsymbol{\hat{\beta}_{1}}$2485.8 3434.2 -2485.8 -5679.01 -578.6 3434.2
$\mathbf{p}$-value $\mathbf{0.0305}$ 0.0001 $\mathbf{0.0305}$ $\mathbf{1.8\times 10^{-5}}$ 0.468 0.0001

### v4

@@ 5.3 Modality
We provide an illustrative example to infer the modality of a smooth regression function from noisy samples using estimated barcodes of its first derivative. We characterize the modes of a smooth regression function $f$ in [-1, 1] using the critical points of $f$. We say that a critical point $c\in$ [-1, 1] is a mode of $f$ if there exists a $\delta>0$ such that $f$ is non-decreasing on $c-\delta,c]$, and non-increasing on $c,c+\delta]$.
We simulate noisy observations from the mixture of two normal distributions truncated on [-1, 1] with means -0.5 and 0.5, with the same variance 0.01, and equal mixture proportions. We choose the kernel bandwidth $h=0.3$ and estimate the regression function for $n=200$. The graph of the true and estimated function using the truncated Gaussian kernel is given in Figure 10 along with the SiZer map of the data. We obtain critical points of the estimated function numerically, using the bisection method. We focus on two of the critical points depicted by green dots in Figure 10, and aim to infer the modality of the underlying true function in the local regions highlighted by grey vertical dotted lines in Figure 10. These regions are $\delta$-neighbourhood around the critical points of the estimated function, where we choose $\delta=0.1$. We use the aforementioned characterization to infer whether the critical points represent modes of the true function.
@@
We obtain barcodes of the estimated derivative using the procedure given in Section 3 and infer the monotonicity of the function in the local regions $c-\delta,c]$ and $c,c+\delta]$ to infer the modality. Note that, here $c$ represents the critical points of interest, and $\delta=0.1$. We estimate barcodes of $\hat{m}_{1}$ in these local regions using the estimated critical points. The barcode of $\hat{m}_{1}$ are obtained using the filtration parameter $\epsilon=0.005$ for the truncated Gaussian and truncated Cauchy kernel on [-1, 1] using $h=0.3$ and $n=200$. We denote the critical points of the estimated function $\hat{m}$ depicted in Figure 10 by $\boldsymbol{\hat{c}_{1}}$ and $\boldsymbol{\hat{c}_{2}}$. The true and estimated barcode are shown in Figure 11 and results are articulated in Table 1.
Table 1: Results for the truncated Gaussian and Cauchy kernel
Estimated critical points ($\hat{c}$) Barcode on $\hat{c}-0.1,\hat{c}$] Barcode on $\hat{c},\hat{c}+0.1]$
$\mathbf{\hat{c}_{1}}$ $\mathbf{\hat{c}_{2}}$ $\hat{c}_{1}-0.1,\hat{c}_{1}] [$$\hat{c}_{2}-0.1,\hat{c}_{2}] [$$\hat{c}_{1},\hat{c}_{1}+0.1] [$$\hat{c}_{2},\hat{c}_{2}+0.1]$
Gaussian -0.54 0.56 [0.05, 0.34] [0.10, 0.41] [-0.43, -0.04] [-0.39, -0.07]
Cauchy -0.51 0.52 [0.07, 0.52] [0.06, 0.71] [-0.63, -0.05] [-0.68, -0.36]
Figure 11: Estimated barcodes of $\hat{m}_{1}$ in local regions $\hat{c}_{i}-0.1,\hat{c}_{i}]$ and $\hat{c}_{i},\hat{c}_{i}+0.1]$, i = 1, 2.
Note that we highlight the true and estimated death or birth times in different regions using green dotted lines in Figure 11 to investigate monotonicity in these regions (see, Section 5.1). The observed barcodes in the local regions suggest that the estimated function is non-decreasing in $\hat{c}_{i}-0.1,\hat{c}_{i}]$ and non-increasing in $\hat{c}_{i},\hat{c}_{i}+0.1]$, $i=1,2$. This is evident from Table 1 and Figure 11 as the observed death times in $\hat{c}_{i}-0.1,\hat{c}_{i}]$ are positive and the observed birth times in $\hat{c}_{i},\hat{c}_{i}+0.1]$ are negative, $i=1,2$. Next, we measure the statistical significance of the observed barcodes by Equation (4.3) using $z_{n,\alpha}=\epsilon$ which is obtained by Equation (4.2) for $\alpha=0.05,h=0.3,\epsilon=0.005,\text{ and }n=200$. Note that, here Equation (4.2) yields $z_{n,\alpha}=\epsilon$ as the constant $C_{\epsilon}$ is large enough, making the estimated barcodes consistent for the aforementioned parameters and sample size. Therefore, as the length of all the observed bars is larger than 0.03, we conclude that the observed bars for both kernels are significant at $\alpha=0.05$ level of significance.
Hence, we conclude that the observed modes are significant at the 95% level of confidence as per the proposed characterization using barcodes of $\hat{m}_{1}$. The SiZer map of the data shown in Figure 10 also suggests the same. The estimated critical points (green dots in Figure 10) lie in a purple region, and the regions on the left and right of these points are colored blue and red, respectively. This suggests the presence of two significant modes in the data for $h=0.3$, which is highlighted by black dotted lines on the SiZer map in Figure 10.
@@ 6 Real data study
We implement the proposed methodology on two real data sets to infer local structures such as monotonicity, convexity, and modality. We measure the statistical significance of observed structures using the methodology proposed in Section 4, and compare the results with the respective SiZer maps. Data set 1 is used to implement the monotonicity characterization using the choice of kernel bandwidth obtained from Equation (4.4) to infer the bars at the 5% level of significance. Data set 2 is used to implement the characterization of convexity and modality using a particular choice of bandwidth and to provide inference using the proposed methodology.
@@ 6.1 Data set 1
We consider the cars data available in the R software library, “datasets.” This data set consists of $n=50$ observations on the speed of cars (X) and the stopping distances of cars (Y). We fit a non-parametric regression function to the observed data using the kernel bandwidth obtained from Equation (4.4) for $\epsilon=0.001$ and $\alpha=0.05$. Note that, to solve Equation (4.4) for $h$, we use estimates of $\gamma_{1}$ (see Assumption (B.5)) and $N_{\epsilon}$ (see Equation 3.6) obtained from Algorithm 1. However, in this case $\eta=Median\{X_{(2)}-X_{(1)},\ldots,X_{(n)}-X_{(n-1)}\}=0$ in Step 2 of Algorithm 1, which should be >> 0. Therefore, we make a slight change in Step 2 (Threshold Selection) of Algorithm 1, and define $\eta=mean\{X_{(2)}-X_{(1)},\ldots,X_{(n)}-X_{(n-1)}\}$ which is >> 0. We refer to Remark 4.1 for more details on the choice of $\eta$ in Algorithm 1. The obtained kernel bandwidth is $h=15.5$.
We fit a non-parametric regression function ($\hat{m}$) to the data, and obtain barcodes of its derivative ($\hat{m}_{1}$) using the procedure given in Section 3. The estimated barcode of $\hat{m}_{1}$ is $\{[\boldsymbol{0.398,0.414}]\}$ which is significant at 5% level of significance by Equation (4.3) for $z_{n,\alpha}=0.001$. This is because the length of the observed bar is larger than 0.007, and the choice of $z_{n,\alpha}$ comes from the fact that $h=15.5$ is a solution of Equation (4.4) for $\epsilon=0.001$ and $\alpha=0.05$. Further, the observed death time is positive, which suggests that the true function underlying the data is strictly increasing. This conclusion is also valid from the SiZer map of the data as shown in Figure 12. The gray dotted line on the SiZer map in Figure 12 highlights $h=15.5$.
Figure 12: The observed data, estimated function for $h=15.5$, SiZer map, and the estimated barcodes of $\hat{m}_{1}$.
@@ 6.2 Data set 2
We consider the “Motorcycle data” available in the R package “adlift”. This data set contains $n=133$ observations of simulations measuring the effects of motorcycle crashes on victims heads. There are two variables in the data set; time after a simulated impact with motorcycles (X) and head acceleration of a post-mortem human test object (Y). The observed structure in the data set is shown in Figure 13. We fit a non-parametric regression model (see, Equation 3.1) to the observed data, and aim to infer the local structures present in the data. In particular, we are interested in inferring the observed convexity structure arising in the local region [15, 25], and modality in [25, 35]. These local regions are highlighted by green dotted lines in Figure 13. A non-parametric regression fit ($\hat{m}$), the estimated derivative ($\hat{m}_{1}$), the SiZer map of the data, and the estimated barcodes of $\hat{m}_{1}$ are shown in Figure 13. The black dotted line on the SiZer map highlights the kernel bandwidth $h=2.5$.
@@
Figure 13: Comparison of local structures in the estimated regression function ($\hat{m}$) derived from the estimated derivative ($\hat{m}_{1}$) with the SiZer map using the bandwidth $h=2.5$
We compute the barcode of the estimated derivative using the procedure given in Section 3 to investigate whether the aforementioned observed local structures are significant or not. We estimate barcodes for the filtration parameter $\epsilon=0.1$ and the kernel bandwidth $h=2.5$. We measure the statistical significance of the observed barcodes using Equation (4.3) for $z_{n.\alpha}$ obtained from Equation (4.2) for $\alpha=0.05$ and for the aforementioned parameters.
The estimated barcode of $\hat{m}_{1}$ on [15, 25] is {[-10, 12.8], [-9.8, -8]}. Now, by Equation (4.2), we obtain $z_{n,\alpha}=0.1$, implying that both observed bars are significant, because the lengths of these bars are larger than 0.7. Therefore, both observed bars are significant at the 95% level of confidence by Equation (4.3). Hence, by step (i) of the characterization of convexity proposed in Section 5.2, we infer that the observed valley structure in [15, 25] is not statistically significant. This conclusion is consistent with the SiZer map in Figure 13. Note that this may not be obvious from the SiZer map of the data shown in Figure 13. However, the SiZer map on [15, 25] in Figure 13 clearly indicates that the observed valley structure is not significant. Because it does not detect the increasing slope with confidence in the local region [15, 25] for $h=2.5$. Moreover, the second bar [-9.8, -8] infers the significant decreasingness of the estimated function in the respective local region, which is around [15, 20]. This conclusion is also consistent with the SiZer map, as it highlights the region [15, 20] in red color in Figure 13.
Next, we estimate the barcode of $\hat{m}_{1}$ on [25, 35] to investigate the peak structure in the region. The estimated barcode is {[-3, 13]}, which is not significant at 5% level of significance using the significance measure given in Equation (4.3). Note that there must be a solution of Equation (4.2) for $z_{n,\alpha}$ in $(\epsilon,16/5\sqrt{2})$ for the observed bar to be significant, as the length of the observed bar is 16. Therefore, since there does not exist any solution of Equation (4.2) in $(\epsilon,16/5\sqrt{2})$ for $z_{n,\alpha}$, we conclude that the observed bar is not significant at the 5% level of significance. Hence, the observed mode in [25, 35] is not significant. The SiZer map in Figure 13 also suggests the same, as it does not detect the falling slope with confidence in the region [25, 35].